find an equation of a parabola whose vertex is (-4,6) and has a solution point of (1,81).
Thanks for the help in advance really need it ha.

Answer by scott8148(6628)

the general vertex form is ___ y = a(x - h)^2 + k , where (h,k) is the vertex

to find a , plug in the given point ___ 81 = a(1 - -4)^2 + 6

75 = a(5)^2 ___ 3 = a

y = 3(x + 4)^2 + 6

more commonly ___ y = 3x^2 + 24x + 54